• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name B. Hahn

    Log entry time 08:20:42 on March26,2007

    Entry number 197538

    keyword=Owl Shift summary

    Kin 5 Bi target need 0.72 C (We have now 0.458)
    ---------------------------------------------------
    Run 1999 40uA, 1M events, 0.045 C
    Run 2000 40uA, 1M events, 0.044 C
    Run 2001 40uA, 1M events, 0.045 C

    Target Ladder Camera is not working. We tried to turn it off and on with no results.

    Run 2002 40uA, 1M events, 0.047 C
    Run 2003 40uA, 1M events, 0.048 C
    Run 2004 40uA, 1M 0.047C

    2:00 am Moved to empty target for harp scan
    2:26 am Harp scan result: Sigma Y < 0.3 --> Asked for Sigma Y < 0.14
    2:50 am After adjustments Harp Scan is fine (Sigma Y = 0.10)
    2:56 am Rebooting Frozen IOC's (They were frozen during harp scan)
    3:25 am Called RC
    3:31 am Asked for Control Access to Hall A to reboot (Cannot connect to hareboot27 to reboot Right NMR)

    4:30 am Moving Spectrometer to KIN 7 (Right arm --> 36.76 deg)
    Floor Mark (37.0 deg) Vernier (mm) -75.8

    04:44:52 Run 2005 Bi Kin07 10uA 0.0059C
    04:56:24 Run 2006 Bi Kin07 20uA 0.02788C
    05:20:56 Run 2007 Bi Kin07 40uA 0.04379C
    05:42:18 Run 2008 Bi Kin07 40uA 0.04268C
    06:02:39 Run 2009 Bi Kin07 40uA 0.04512C
    06:34:14 Run 2010 Bi Kin07 40uA 0.04362C
    06:54:50 Run 2011 Bi Kin07 40uA 0.04402C
    07:16:37 Run 2012 Bi Kin07 40uA 0.04372C
    07:36:04 Run 2013 Bi Kin07 40uA 0.04473C
    07:57:37 Run 2014 Bi Kin07 40uA 0.008213